Butane gas, C4H10, is combusted to produce carbon dioxide and water vapour. If 0.67 mol of butane are reacted, how many moles of water vapour are produced?
1
Expert's answer
2020-10-26T14:54:42-0400

2C4 H10 + 13 O2 = 8 CO2 +10 H2O

n (C4 H10 )= 0,67 mole  
n (H2O)=10*0.67/2=3,35 mole


answer: 3,35 mole